OpenStudy (anonymous):

Given the equation 4x − 26 = 2x − 4, which order of operations completely solves for x? Subtract 2x, then add 26, lastly divide by 2 Subtract 4x, then add 26, lastly divide by 6 Subtract 4x, then subtract 4, lastly divide by 6 Subtract 2x, then subtract 4, lastly divide by 6

OpenStudy (hanner_b_nanner):

ok, how do we start to solve for this? Do you know? :)

OpenStudy (anonymous):

lol idk

OpenStudy (hanner_b_nanner):

well, you have to subtract the lowest number of x from the higher number. 4x − 26 = 2x − 4 -2x -2x __________________ 2x-26= -4 Then, you just finish it like a regular equation. 2x-26= -4 +26 +26 _______________ 2x=22 _______ 2 x=2 So, the answer will be...A!!! :)))

OpenStudy (anonymous):

ohhh thank yu (-:

OpenStudy (hanner_b_nanner):

anyyytimmeee! :D